OpenStudy (anonymous):

Consider this equation: -3x - 5 + 2x = 6 Which is an equivalent equation after combining like terms? A. -8 + 2x = 6 B. -3x - 3 = 6 C. -3x - 5 = 4 D. -x - 5 = 6

OpenStudy (anonymous):

@FibonacciChick666 please help

OpenStudy (fibonaccichick666):

well, what are your like terms?

OpenStudy (anonymous):

-3x and 2x but idk what to do after that and what the answer would be

OpenStudy (fibonaccichick666):

well, what do you get when you combine those?

OpenStudy (fibonaccichick666):

ie, -3x+2x=?

OpenStudy (fibonaccichick666):

-3x-5+2x=6----> (-3x+2x)-5=6. I have to go shower, but use your order of ops and you'll get it

OpenStudy (anonymous):

its either -2x or its 5x

OpenStudy (fibonaccichick666):

-3+2?

OpenStudy (fibonaccichick666):

draw it on a number line if you don't know for sure

OpenStudy (anonymous):

i meant -1 not -2

OpenStudy (fibonaccichick666):

(it is neither of those)

OpenStudy (fibonaccichick666):

yes, that's right

OpenStudy (fibonaccichick666):

so then, plug and chug (-3x+2x)-5=6.

OpenStudy (fibonaccichick666):

well simplify the parentheses

OpenStudy (anonymous):

wouldn't it be D

OpenStudy (fibonaccichick666):

yup

OpenStudy (anonymous):

thanks so much

OpenStudy (fibonaccichick666):

np. Just remember to rearrange and put in parentheses to make it more obvious

OpenStudy (anonymous):

okayy